How much 85% acid solution must be added to 50ml of a 25% acid solution to produce a mixture that is at 60%acid?

Answer by stanbon(75887)

How much 85% acid solution must be added to 50ml of a 25% acid solution to produce a mixture that is at 60%acid?
======
acid + acid = acid
0.25*50 + 0.85x = 0.60(50+x)
--------
25*50 + 85x = 60*50 + 60x
25x = 35*50
x = 70 ml (amt of 85% solution needed)
----------------
Cheers,
Stan H.